OpenStudy (anonymous):

Find any vertical asymptots f(x) = 1 / |x-4|

OpenStudy (across):

Vertical asymptotes are points where the slope of a curve blows up to infinity, i.e., it's undefined. You have a fraction there. What's so peculiar about fractions with variables in the denominator?

OpenStudy (anonymous):

i think the teacher wants us to sketch this function.

OpenStudy (across):

Graph it point-wise or by inspection? Also, hint: what would happen if x = 4?

OpenStudy (anonymous):

i don't know, that's why i am posting this question.

OpenStudy (across):

\[f(x)=\frac{1}{|x-4|}=\frac{1}{|(4)-4|}=\frac{1}{0}?\]

OpenStudy (across):

Can you divide by zero?

OpenStudy (anonymous):

no

OpenStudy (across):

There's your asymptote.

OpenStudy (anonymous):

i think its a sketch , it would be like this..